# FARM: Functional Group-Aware Representations for Small Molecules

### v3

@@ Abstract
We introduce Functional Group-Aware Representations for Small Molecules (FARM), a novel foundation model designed to bridge the gap between SMILES, natural language, and molecular graphs. The key innovation of FARM lies in its functional group-aware tokenization, which directly incorporates functional group information into SMILES, enriching SMILES with detailed chemical context. For example, instead of using “O” to represent all oxygen atoms, we use specific tokens like “O_ketone” and “O_hydroxyl” to differentiate oxygen atoms belonging to distinct functional groups. This tokenization expands the chemical lexicon, effectively bridging the gap between SMILES and natural language in terms of vocabulary size, ultimately enhancing the model’s ability to predict molecular properties. FARM also represents molecules from two perspectives: by (1) using masked language modeling to capture atom-level features and (2) employing graph neural networks to encode the whole molecule topology. FARM leverages contrastive learning to aligns these two views of representations into a unified molecular embedding. We rigorously evaluate FARM on the MoleculeNet dataset, where it achieves state-of-the-art performance on 11 out of 13 tasks. These results highlight FARM’s potential to improve molecular representation learning and demonstrate its strong transfer learning capabilities, paving the way for promising applications in drug discovery and pharmaceutical research. The code is available at: https://github.com/thaonguyen217/farm_molecular_representation.
@@ 1 Introduction
Artificial intelligence (AI) has emerged as a transformative tool in accelerating scientific discovery, particularly in drug development. It is increasingly employed for tasks such as molecular property prediction, drug-target interaction prediction, and quantitative structure-activity relationship (QSAR) modeling (Chen et al., 2016; Wen et al., 2017; Shen and Nicolaou, 2019; Walters and Barzilay, 2020; Achary, 2020; Wang et al., 2022a; Edwards et al., 2022; Zhang et al., 2023a; Nguyen et al., 2024a; Edwards et al., 2024b, a). However, one of the central challenges in this field is the scarcity of large labeled datasets required for traditional supervised learning methods. This has shifted the focus towards self-supervised pre-trained models that can extract meaningful patterns from vast amounts of unlabeled molecular data (Shen and Nicolaou, 2019). As a result, the development of robust foundation models for molecular representations is now more critical than ever.
Despite significant advancements in other domains, such as natural language processing (NLP) and computer vision, there is still no dominant foundation model tailored to molecular representation in drug discovery (Zhang et al., 2023b). This paper begins to address this pressing gap by introducing an innovative approach that leverages functional group (FG)-aware tokenization in the context of both sequence-based and graph-based molecular representations. In the literature, terms such as “motifs,” “fragments,” “substructures,” and “building blocks” are often used interchangeably with “functional groups.” However, in this paper, we use the term “functional groups” to represent a subset of these molecular concepts, with a more rigorous definition grounded in chemical principles. A functional group refers to a chemically meaningful portion of a molecule that significantly influences its properties and behavior. This can include simple functional groups, such as hydroxyl (-OH), as well as more complex molecular substructures, such as ring systems, which serve distinct functional roles within the molecule. Functional groups (FGs) play a crucial role in determining a molecule’s properties. For instance, salicylic acid and aspirin share the same core structure but differ by just one functional group—a hydroxyl group in salicylic acid is replaced by an acetyl group in aspirin. This seemingly minor change results in substantial differences in their chemical properties and biological activities. The acetyl group in aspirin enables it to irreversibly inhibit cyclooxygenase (COX) enzymes, providing it with blood-thinning properties that make it effective in preventing heart attacks and strokes—functions that salicylic acid does not exhibit.
The innovation here lies in expanding tokenization granularity for SMILES and reducing tokenization granularity for molecular graphs in a way that explicitly interfaces with key drivers of molecular function. This approach enables molecular embeddings to more effectively support models in understanding and predicting molecular functions. For example, instead of representing all oxygen atoms with the same token (e.g., “O”), we use more specific tokens such as “O_ketone” and “O_hydroxyl” to distinguish oxygen atoms that belong to different functional groups in the SMILES string. This approach addresses a major limitation of sequence-based models, which typically focus only on individual atom types while neglecting the higher-level functional groups crucial for accurate molecular understanding. FG-aware tokenization enriches the SMILES representation with chemically relevant context, bridging the gap between the expansive vocabulary of natural language—the language for which Transformers were originally designed—and the limited chemical lexicon of the SMILES language, thereby reducing negative transfer during learning.
Conversely, instead of representing a ketone group (C=O) using two separate nodes in a graph, we represent it as a single node corresponding to the functional group as a whole. This abstraction yields a high-level “skeleton” of the molecule, where each node denotes a chemically meaningful functional unit rather than individual atoms. To capture interactions between these units, we construct a functional group (FG) knowledge graph that encodes chemically informed embeddings for each functional group based on their properties and co-occurrence patterns. These embeddings are then leveraged in a link prediction task to learn structural representations of molecules—allowing the model to infer both functional group similarity and likely connectivity. This FG graph captures structural relationships that are often implicit or obscured in SMILES representations. As a result, the functional group graph provides a complementary perspective to the FG-enhanced SMILES, encoding topological and structural context, while the enhanced SMILES preserves fine-grained sequence-level and chemical detail. Contrastive learning will then be used to align these two views, encouraging the model to learn representations that are sensitive to chemically meaningful structures. Figure 1 and Figure 2a present an overview of our molecular representation learning model.
@@
  • FG-aware tokenization and fragmentation: We introduce FG-aware tokenization and fragmentation, adding rich chemical context to each atom and bridging the gap between SMILES and natural language.
  • Structural representation learning: We leverage a FG knowledge graph for robust FG embeddings and effectively learn molecular structure through link prediction between FGs.
  • Atom-feature and structural representation integration: We combine masked language model for learning atom-level features with GNNs for capturing structural information.
  • Robustness in downstream tasks: FARM demonstrates strong transfer learning capabilities - the core goal of pretrained models - outperforming other methods in 11 out of 13 tasks from the MoleculeNet benchmark.
@@ 2 Related Work
Functional group-Aware Molecular Representations   In recent years, there has been growing recognition that incorporating functional group (FG) information into molecular representations can significantly enhance model performance in downstream tasks. Approaches in this domain can be broadly classified into two categories: those leveraging language models (LMs) (Li et al., 2023; Xia et al., 2023) and those utilizing graph neural networks (GNNs) (Zhang et al., 2020, 2021; Yu and Gao, 2022; Yang et al., 2022; Wang et al., 2023; Wu et al., 2023; Han et al., 2023; Fang et al., 2023). Within these categories, methods either employ rule-based functional group detection, relying on predefined chemical rules to identify FGs, or adopt unsupervised strategies that infer substructures or motifs from the data. Regardless of the approach, these models consistently demonstrate that enriching molecular representations with functional group information leads to improved performance across a wide range of molecular property prediction tasks (Fang et al., 2023; Han et al., 2023; Wang et al., 2023). This underlines the importance of functional group awareness in achieving accurate and generalizable molecular representations.
@@
In studies such as Zhang et al. (2021); Han et al. (2023); Chen et al. (2024); Yang et al. (2022), the BRICS algorithm (Degen et al., 2008) is employed to fragment molecules, with some extending the approach through additional rules to achieve finer-grained segmentation. However, this method focuses on general reaction-based bond breaking, rather than targeting functional group-specific characteristics. Additionally, while BRICS groups atoms into fragments, it does not explicitly identify or label functional groups, limiting its utility in tasks that require functional groups-specific information. Other works (Li et al., 2023; Chen et al., 2024) leverage RDKit (Landrum, 2010) for functional group detection. While RDKit is effective in identifying common functional groups, its capabilities are limited to well-known groups and do not extend to detecting more complex functional groups, such as ring systems that often dominate molecular datasets and play a critical role in molecular function. In this work, we present a rule-based functional group detection algorithm that accurately identifies 101 common groups and all ring-containing ones. Unlike frequent subgraph mining or motif-based methods, our approach is explicitly grounded in chemical principles, ensuring reliable and chemically meaningful detection aligned with expert definitions.
@@
Contrastive Learning-Based Molecular Representations   Contrastive learning, a self-supervised learning technique, aims to learn representations by maximizing agreement between positive pairs while distinguishing them from negative samples. In the context of molecular representation, Wang et al. (2022b) applies this technique by augmenting each molecular graph to create slightly different versions, which are then treated as negative examples to enhance the learning process. In Pinheiro et al. (2022), the authors use a molecular graph encoder and a SMILES encoder to encode both molecular graphs and SMILES, employing contrastive loss to maximize the agreement between these embeddings. This method enriches SMILES with topology information and the molecular graph with sequence context, making the final embeddings of molecules more robust. In Zhang et al. (2020), the authors minimize the distance between the representation of a molecule and the representations of its constituent substructures. Collectively, these works highlight the versatility of contrastive learning in unifying diverse molecular representations, leading to improved downstream performance in molecular tasks. Our method extends this approach by using contrastive learning to maximize the agreement between FG-enhanced SMILES representations and FG graphs, thereby enhancing the alignment between sequence-based and graph-based representations, ensuring a more comprehensive and chemically informed embedding.
@@ 3 Methodology
We present our methodologies for enhancing molecular representation learning. Section 3.1 introduces FG-aware tokenization and fragmentation, which injects detailed chemical context into both sequence and graph representations. Section 3.2 delves into the masked atom prediction task, a self-supervised technique that enables the model to learn atom-level representations. In Section 3.3, we describe our method for capturing the core molecular structure. Finally, Section 3.4 presents contrastive learning, which aligns FG-enhanced SMILES strings with FG graph embeddings to achieve a comprehensive, unified molecular representation.
@@ 3.1 Functional Group-Aware Tokenization and Fragmentation
We propose an FG-aware tokenization and fragmentation method that embeds detailed functional group information into molecular representations, tailoring it for both SMILES and graph-based models. This method defines a set of functional groups and employs an algorithm to detect them within the molecular graph. The algorithm traverses the graph, evaluating each atom based on criteria such as atom type, neighboring atom types and corresponding bonds, number of neighbors, atom charge, and bonded hydrogen atoms to identify the functional group. This approach ensures precise detection of functional groups within molecular graphs. For instance, a carbon atom with a charge of 0 and three neighbors, including a double-bonded oxygen atom, is classified as part of a ketone group (RCOR’), with the oxygen also contributing to the group. Additionally, we address cases where a functional group may be a subset of another. The algorithm first checks for the presence of the larger functional group. If it’s not identified, the algorithm then checks for smaller functional groups, ensuring correct identification even in complex structures. Appendix B.1 provides a full list of 101 non-ring-containing functional groups.
The results of the FG detection algorithm are utilized for two distinct processes:
  • FG-aware tokenization: The molecular graph, where each node (atom) is assigned to a specific functional group, is converted back into a SMILES string that incorporates functional group information. For example, a ketone-containing group originally depicted in SMILES as *C(=O)* is transformed into an FG-enhanced SMILES string like *C_ketone(=O_ketone)*. This FG-enhanced SMILES embeds additional chemical context directly into the molecular representation while remaining fully compliant with traditional SMILES rules (if the FG information is removed, the FG-enhanced SMILES reverts to its standard SMILES form). Similar to how words like “bank” have multiple meanings depending on context (e.g., river bank vs. financial bank), our method captures fine-grained semantic distinctions by representing functional groups with more precise tokens—for example, distinguishing ketone oxygen from hydroxyl oxygen.
  • FG-aware fragmentation: Once functional groups are identified, the molecule is segmented based on the bonds connecting these groups, as illustrated in Figure 2(b). These bonds are represented as edges in a graph, where the nodes correspond to the functional groups. This structure, known as the functional group graph, conveys the molecule’s structural information.

### v4

# FARM: Enhancing Molecular Representations with Functional Group Awareness

@@ Abstract
We introduce Functional Group-Aware Representations for Small Molecules (FARM), a novel foundation model designed to bridge the gap between SMILES (a linear string representation of molecular structures), natural language, and molecular graphs. The key innovation of FARM lies in its functional group (FG) annotation at the atomic level, which enables both FG-enhanced SMILES and FG graphs: SMILES are enriched with FG information to specify which functional group each atom belongs to, while the FG graph captures the molecular backbone by showing how the functional groups are connected. This tokenization not only injects chemical knowledge into SMILES but also expands the chemical lexicon, effectively bridging the gap between SMILES and natural language in terms of vocabulary size, making the sequences more suitable for Transformer-based models. FARM then learns molecules from two complementary perspectives to fully encode both functional and structural information. Masked language modeling on FG-enhanced SMILES captures atom-level features enriched with FG context, while graph neural networks encode higher-level molecular topology by representing how FGs are connected. Contrastive learning then aligns these two views into a unified embedding, ensuring that atom-level details and FG-level structure are jointly represented. This dual-level modeling is central to FARM’s ability to predict molecular properties accurately. We rigorously evaluate FARM on the MoleculeNet dataset, achieving state-of-the-art performance on 11 out of 13 tasks, and further validate its generalization on the photostability dataset for quantum mechanics properties. These results highlight FARM’s potential to improve molecular representation learning, demonstrate its strong transfer learning capabilities across drug discovery and material design domains, and make way for broad applications in pharmaceutical research and functional materials development. The code is available at: https://anonymous.4open.science/r/farm-E3CB.
@@ 1 Introduction
Artificial intelligence (AI) has emerged as a transformative tool in accelerating scientific discovery, particularly in drug development and material design. In these domains, it is increasingly employed for tasks such as molecular property prediction, drug–target interaction prediction, quantitative structure–activity relationship (QSAR) modeling, and the discovery of novel materials with tailored electronic, mechanical, or optical properties. (Chen et al., 2016; Wen et al., 2017; Shen and Nicolaou, 2019; Walters and Barzilay, 2020; Achary, 2020; Wang et al., 2022a; Edwards et al., 2022; Zhang et al., 2023a; Nguyen et al., 2024a; Edwards et al., 2024b; a). However, one of the central challenges in this field is the scarcity of large labeled datasets required for traditional supervised learning methods. This has shifted the focus towards self-supervised pre-trained models that can extract meaningful patterns from vast amounts of unlabeled molecular data (Shen and Nicolaou, 2019). As a result, the development of robust foundation models for molecular representations is now more critical than ever.
Despite significant advancements in other domains, such as natural language processing (NLP) and computer vision, there is still no dominant foundation model tailored to molecular representation in drug discovery (Zhang et al., 2023b). This paper begins to address this pressing gap by introducing an innovative functional group (FG)-aware approach, in which classical FGs are identified according to established chemical principles, while also considering additional chemically meaningful substructures that strongly influence molecular properties. Functional groups play a central role in determining the chemical and biological behavior of molecules, and even small changes can lead to significant differences in activity. For instance, salicylic acid and aspirin share the same molecular backbone but differ by a single functional group: the hydroxyl group in salicylic acid is replaced by an acetyl group in aspirin. This subtle modification dramatically alters their properties and biological activities, as the acetyl group in aspirin enables irreversible inhibition of cyclooxygenase (COX) enzymes, producing blood-thinning effects that help prevent heart attacks and strokes, effects not observed with salicylic acid. By rigorously grounding tokenization in chemical rules, our approach aims to more effectively capture these critical molecular properties and functions.
In the strictest chemical sense, functional groups refer to well-defined moieties such as hydroxyl, carbonyl, or amino groups. In this work, we adopt a broader interpretation of FGs to include not only these classical groups but also ring systems and fused ring systems, which, although not always labeled as functional groups in textbooks, often play critical roles in molecular recognition, stability, and activity.
The key innovation of our approach lies in how molecules are tokenized. In SMILES representations, we enrich detail by annotating each atom with its FG, while in molecular graph representations, we abstract detail to emphasize the molecular backbone, namely which FGs are present and how they are connected. This dual strategy allows molecular embeddings to more faithfully capture functional information, enabling more accurate prediction of molecular properties and functions.
For example, instead of representing every oxygen atom with the same token (“O”), we assign different tokens based on their role in functional groups, such as “O_ketone” or “O_hydroxyl”(Figure 1). Traditional sequence-based models typically treat atoms only by type and ignore the functional groups they belong to, even though a single atom can play very different roles depending on its chemical context. This is analogous to natural language, where a word like “bank” can mean either a river bank or a financial institution. By specifying an atom’s functional group, our approach enriches SMILES with chemically contextual information. We refer to this representation as FG-enhanced SMILES, which enriches molecular sequences with chemically meaningful context and expands the vocabulary beyond the limited set of atom symbols in standard SMILES.
This vocabulary expansion is particularly important because standard SMILES have a very limited and repetitive vocabulary, essentially atom and bond types, with the majority being C, O, and N. Training models on such a restricted corpus often leads to rapid convergence, but the representations capture little about functional or property-related patterns. As a result, downstream performance can suffer and even exhibit negative transfer (Xia et al., 2023). By enriching SMILES with FG annotations, we enlarge the vocabulary and bring it closer in spirit to natural language, the domain for which Transformer models are designed. This richer representation helps bridge the gap between SMILES and natural language, thereby mitigating negative transfer and improving the model’s ability to learn molecular properties and functions.
While FG-enhanced SMILES captures fine-grained chemical details, sequential representations like SMILES often struggle to represent the overall molecular backbone– the higher level arrangement of functional groups that defines molecular topology. To address this, we construct a FG graph, where each node corresponds to a functional group and edges encode their connectivity. For example, a ketone group (C=O) is represented as a single FG node rather than two separate atoms, producing a high-level molecular backbone. Each FG node is assigned a chemically informed embedding, learned from a FG knowledge graph that captures both the intrinsic properties of each FG and their relationships with other FGs across molecules. These embeddings are then used to initialize a graph neural network (GNN), which is trained via a link prediction task to reconstruct connections between FGs. By predicting which FGs are likely to be connected, the GNN learns the molecule’s high-level structure, relationships that are often implicit in SMILES representations.
Finally, contrastive learning is employed to align the two views: the FG-enhanced SMILES and the FG graph, encouraging the model to learn unified representations that are sensitive to chemically meaningful structures. Figure 1 and Figure 2a present an overview of our molecular representation learning model.
@@
  • FG-aware tokenization and fragmentation: We introduce an FG annotation algorithm that labels each atom in a molecule with the FG to which it belongs. The FG-annotated molecular graph is then used for FG-aware tokenization and fragmentation, producing both FG-enhanced SMILES and FG graphs. This approach enriches each atom with chemical context, bridges the gap between SMILES and natural language through a FG–aware vocabulary, and still captures higher-level structural information in the graph.
  • Structural representation learning: We construct a FG knowledge graph to obtain robust FG embeddings, and learn molecular structure by predicting the links between FGs, thereby capturing the high-level structural organization of molecules.
  • Atom-feature and structural representation integration: We combine a masked language model to learn atom-level features with GNNs to capture structural information, and align these two views using contrastive learning.
  • Robustness in downstream tasks: FARM demonstrates strong transfer learning capabilities, the core goal of pretrained models, outperforming other methods in 11 out of 13 tasks from MoleculeNet, which spans physiology, biophysics, physical chemistry, and quantum physics. Beyond MoleculeNet, FARM further shows robust generalization on quantum mechanics datasets, including the photostability dataset.
@@ 2 Related Work
Functional group-Aware Molecular Representations In recent years, there has been growing recognition that incorporating functional group (FG) information into molecular representations can significantly enhance model performance in downstream tasks. Approaches in this domain can be broadly classified into two categories: those leveraging language models (LMs) (Li et al., 2023; Xia et al., 2023) and those utilizing graph neural networks (GNNs) (Zhang et al., 2020; 2021; Yu and Gao, 2022; Yang et al., 2022; Wang et al., 2023; Wu et al., 2023; Han et al., 2023; Fang et al., 2023). Within these categories, methods either employ rule-based functional group detection, relying on predefined chemical rules to identify FGs, or adopt unsupervised strategies that infer substructures or motifs from the data. Regardless of the approach, these models consistently demonstrate that enriching molecular representations with functional group information leads to improved performance across a wide range of molecular property prediction tasks (Fang et al., 2023; Han et al., 2023; Wang et al., 2023). This underlines the importance of functional group awareness in achieving accurate and generalizable molecular representations. More recently, SCAGE (Qiao et al., Technical report) further validates this direction by incorporating functional group prediction as one of four pretraining objectives in a self-conformation-aware graph transformer, demonstrating that explicitly modeling functional groups at the atomic level leads to improved generalization across diverse molecular property prediction tasks.
@@
In studies such as (Zhang et al., 2021; Han et al., 2023; Chen et al., 2024; Yang et al., 2022), the BRICS algorithm (Degen et al., 2008) is employed to fragment molecules, with some extending the approach through additional rules to achieve finer-grained segmentation. However, this method focuses on general reaction-based bond breaking, rather than targeting functional group-specific characteristics. Additionally, while BRICS groups atoms into fragments, it does not explicitly identify or label functional groups, limiting its utility in tasks that require functional groups-specific information. Other works (Li et al., 2023; Chen et al., 2024) leverage RDKit (Landrum, 2010) for functional group detection. While RDKit is effective in identifying common functional groups, its capabilities are limited to well-known groups and do not extend to detecting more complex functional groups, such as ring systems that often dominate molecular datasets and play a critical role in molecular function. In this work, we present a rule-based functional group detection algorithm that accurately identifies 101 common groups and all ring-containing ones. Unlike frequent subgraph mining or motif-based methods, our approach is explicitly grounded in chemical principles, ensuring reliable and chemically meaningful detection aligned with expert definitions.
@@
Contrastive Learning-Based Molecular Representations Contrastive learning, a self-supervised learning technique, aims to learn representations by maximizing agreement between positive pairs while distinguishing them from negative samples. In the context of molecular representation, (Wang et al., 2022b) applies this technique by augmenting each molecular graph to create slightly different versions, which are then treated as negative examples to enhance the learning process. In (Pinheiro et al., 2022), the authors use a molecular graph encoder and a SMILES encoder to encode both molecular graphs and SMILES, employing contrastive loss to maximize the agreement between these embeddings. This method enriches SMILES with topology information and the molecular graph with sequence context, making the final embeddings of molecules more robust. In (Zhang et al., 2020), the authors minimize the distance between the representation of a molecule and the representations of its constituent substructures. Collectively, these works highlight the versatility of contrastive learning in unifying diverse molecular representations, leading to improved downstream performance in molecular tasks. Our method extends this approach by using contrastive learning to maximize the agreement between FG-enhanced SMILES representations and FG graphs, thereby enhancing the alignment between sequence-based and graph-based representations, ensuring a more comprehensive and chemically informed embedding.
Algorithm 1 Functional group annotation for molecular graph
0: Molecular graph $G$, functional group definitions FG_listFG\_list (ordered largest to smallest)
1:for each atom $v$ in $G$ do
2:  if $v$ does not have a functional group assigned then
3:   $FG\leftarrow$ largest functional group in FG_listFG\_list that matches $v$ and its neighborhood
4:   $matched\_atoms\leftarrow$ atoms involved in the match
5:   for each atom $a$ in matched_atomsmatched\_atoms do
6:    $a$.SetProp("FunctionalGroup", FGFG.name)
7:   end for
8:  end if
9:end for
@@ 3 Methodology
We present our approach for enhancing molecular representation learning through FG-enhanced SMILES learning with Transformers, FG graph backbone modeling, and dual-perspective contrastive learning, jointly capturing atom-level chemical context and FG-level molecular structure. Section 3.1 introduces FG-aware tokenization and fragmentation, which injects detailed chemical context into both sequence and graph representations. Section 3.2 delves into the masked atom prediction task, a self-supervised technique that enables the model to learn atom-level representations. In Section 3.3, we describe our method for capturing the core molecular structure. Finally, Section 3.4 presents contrastive learning, which aligns FG-enhanced SMILES strings with FG graph embeddings to achieve a comprehensive, unified molecular representation.
@@ 3.1 Functional Group-Aware Tokenization and Fragmentation
We propose an FG-aware tokenization and fragmentation method that incorporates detailed functional group information into molecular representations, designed for both SMILES and graph-based models. Our approach defines a set of functional groups and uses an algorithm to detect them within the molecular graph, annotating each atom with its corresponding FG. The algorithm traverses the molecular graph, evaluating each atom based on criteria such as atom type, neighboring atom types and bonds, number of neighbors, atom charge, and bonded hydrogen atoms to identify its functional group. All input molecules are canonicalized using RDKit prior to running this algorithm, ensuring a deterministic atom ordering. This approach ensures precise detection that strictly aligns with the chemical definitions of functional groups within molecular graphs. The pseudocode is provided in Algorithm 1. For instance, a carbon atom with a charge of 0 and three neighbors, including a double-bonded oxygen atom, is classified as part of a ketone group (RCOR’), with the oxygen also contributing to the group. Additionally, we address cases where a functional group may be a subset of another. The algorithm first checks for the presence of the larger functional group. If it’s not identified, the algorithm then checks for smaller functional groups, ensuring correct identification even in complex structures. Appendix B.1 provides a full list of 101 non-ring-containing functional groups.
The results of the FG annotation algorithm are utilized for two distinct processes:
  • FG-aware tokenization: The molecular graph, where each node (atom) is assigned to a specific functional group, is converted back into a SMILES string that incorporates functional group information. For example, a ketone-containing group originally depicted in SMILES as *C(=O)* is transformed into an FG-enhanced SMILES string like *C_ketone(=O_ketone)*. This FG-enhanced SMILES embeds additional chemical context directly into the molecular representation while remaining fully compliant with traditional SMILES rules (if the FG information is removed, the FG-enhanced SMILES reverts to its standard SMILES form).
  • FG-aware fragmentation: Once FGs are identified, the molecule is segmented according to the bonds connecting these groups, as illustrated in Figure 2(b). These bonds serve as edges in a graph, with nodes corresponding to the functional groups. This representation, referred to as the FG graph, captures the molecule’s structural backbone.